mixture a is 30 percent suger, 15 percent ciiamon, and the rest flour. Mixture b is 3/4 sugar and 1/4 flour. If we comine 2 cups of mixture a with 5 cups of mixture b and 1 cup of flour, what percent of the resluting mixture will be flour?

Answer by lwsshak3(11628)

mixture a is 30 percent suger, 15 percent ciiamon, and the rest flour. Mixture b is 3/4 sugar and 1/4 flour. If we comine 2 cups of mixture a with 5 cups of mixture b and 1 cup of flour, what percent of the resulting mixture will be flour?
**
let x=% flour in resulting mixture
..
mixture a: 2 cups
sugar=30%*2=0.6 cup
cinnamon=15%*2=0.3 cup
flour=55%*2=1.1 cup
..
mixture b: 5 cups
sugar=3/4*5=15/4 cup
flour=1/4*5=5/4=1.25 cup
..
Resulting mixture: 8 cups
flour,x=1.1+1.25+1=3.35 cups
x=3.35/8≈.419
% flour in resulting mixture≈41.9%